OpenStudy (anonymous):

solve the equation over the interval [0 degrees, 360 degrees] tan^2theta+5tan theta+3=0 type answer in degrees

OpenStudy (mathmath333):

\(\large\tt \color{black}{tan^2\theta+5tan\theta+3=0}\) \(\large \large\tt \color{black}{put~tan\theta=x}\) \(\large\tt \color{black}{x^2+5x+3=0}\)

OpenStudy (anonymous):

thank you. But the answer should be in degrees, solution set. which is why I am so lost :(

OpenStudy (mathmath333):

u will get value of x in decimals and by further applying 'arctan' u will get theta in degrees

OpenStudy (mathmath333):

\(\large \tt \color{black}{\text{use formula for }}\) \[ax^2+bx+c=0\] \(\large \tt \color{black}{x=\dfrac{-b\pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (mathmath333):

after getting x just u have to use' arctan' on bothsides